OpenStudy (anonymous):

integrate e^x(2^x-x)dx

OpenStudy (anonymous):

\[\int\limits_{}^{}e^x(2^x-x)\]

OpenStudy (anonymous):

Oh wow, holy crap.

OpenStudy (anonymous):

yea that's what I said

OpenStudy (anonymous):

Remember that \[ 2 = e^{\ln2} \implies 2^x = e^{x\ln 2} \]

OpenStudy (anonymous):

Use distributive property.

OpenStudy (anonymous):

then just do each integral separately.

OpenStudy (anonymous):

You can use integration by parts. \\[\int\limits\limits_{}^{} a^x = a^x \ln (a)\ \]

OpenStudy (anonymous):

can you help me with that?

OpenStudy (anonymous):

i mean \[e^x(e^{xln2}-x)\]

OpenStudy (anonymous):

use distributive property first

OpenStudy (anonymous):

meaning?

OpenStudy (anonymous):

multiply out the brackets :P

OpenStudy (anonymous):

\[ c(a+b) = ca+cb \]

OpenStudy (anonymous):

oh yes, of course

OpenStudy (anonymous):

\[ \int f(x)+g(x)dx = \int f(x)dx + \int g(x)dx \]

OpenStudy (anonymous):

\( xe^x \) requires integration by parts. The other one is just chain rule.

OpenStudy (anonymous):

I mean u sub.

OpenStudy (anonymous):

xe^x requires u sub right?

OpenStudy (anonymous):

No it requires integration by parts. You wanna do \(u=x\) and \(dv = e^x\)

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

\[ (uv)' = u'v + uv' \implies \int udv = uv - \int vdu \]

OpenStudy (anonymous):

aren't the both xe^x if you pull that 2 out in front of the integratl sign?

OpenStudy (anonymous):

yes i have memorized that formula

OpenStudy (anonymous):

No one is just \(e^{x(\ln(2) + 1)}\) the other is \(xe^x\)